# Review: Quantum Metrology and Sensing with Many-Body Systems

### v1

@@ II.4 Multi-parameter estimation
and $I(\boldsymbol{\theta})$ is the classical Fisher information matrix with elements Liu et al. (2019)
@@
One of the main differences between the single- and multi-parameter scenarios is that, while the classical Fisher information can be achieved in both scenarios, in principle, the QFI is attainable only in the single-parameter case and not for the jointly multi-parameter estimation Albarelli et al. (2020). This can be understood because, generally, there is no single symmetric logarithmic derivative $L_{i}$ that defines an optimal measurement basis suitable for all the unknown parameters. To quantify this notion of incompatibility between parameters Ragy et al. (2016)—those that cannot be estimated optimally with a single set of measurements—the incompatibility matrix $\boldsymbol{D}$ has been introduced Albarelli et al. (2020); Carollo et al. (2019a), also known as mean Uhlmann curvature Carollo et al. (2019a), with elements:
@@ II.5 Geometry of parameter estimation
Close links between Fisher information quantities and statistical distances have been found Wootters (1981); Braunstein and Caves (1994). In terms of the classical Fisher information, which quantifies the sensitivity of the probability distributions $p(x|\theta)$ concerning the unknown parameter $\theta$, the notion of distinguishability between different probability distributions among $M$ trials arises naturally (see Bengtsson and $\dot{Z}$yczkowski Lehmann and Casella (1998), and Kok and Lovett Kok and Lovett (2010)). It turns out that distinguishing between two probabilities $p(x|\theta_{1})$ and $p(x|\theta_{2})$, taking $\text{d}s$ ($\text{d}s$ being the infinitesimal distance on the probability simplex space) along a line element $\text{d}\theta$, results in Lehmann and Casella (1998); Kok and Lovett (2010)
@@
which is the classical Fisher information of Eq. (3). Further analysis leads to:
@@
where $\delta\theta$ is the segment of the path in the probability simplex—rather than the variance of $\tilde{\theta}$. Note that the above has been derived solely from a distance measure (a metric) by distinguishing probability distributions on the simplex space. It has also been shown that for distinguishing two probability distributions, one can use the relative entropy (not a metric) for this purpose, and consequently, a close connection between the classical Fisher information matrix and the relative entropy emerges Sidhu and Kok (2020).
@@
For pure states $\rho_{1}=|\psi_{1}\rangle\langle\psi_{1}|$ and $\rho_{2}=|\psi_{2}\rangle\langle\psi_{2}|$ the Uhlmann fidelity reduces to $F(\rho_{1},\rho_{2})=|\langle\psi_{1}|\psi_{2}\rangle|$.
@@ II.6 Alternatives to QFI for Metrology
where $I(\theta)$ is the QFI for $\rho(\theta)$ with respect to the parameter $\theta$. One can note that this bound reduces to the biased version of the Bayesian Quantum Cramér-Rao bound bound for a uniform prior. Secondly, and of significant relevance to many-body sensing, the QFI is not fully expressible in terms of experimentally easily accessible one and two-point correlation functions(Gabbrielli et al., 2018). This was addressed recently in Ref. (Lepori et al., 2023), where an alternative functional dependent only on one and two-point correlators was proposed.
@@ IV Equilibrium Quantum Sensing: Second Order Quantum Phase Transition
An alternative method to interferometric sensing is to harness the interaction between the particles in strongly correlated many-body systems. The interplay between different parts of the Hamiltonian can induce phase transition in the system, also known as quantum criticality. Many-body systems exhibiting quantum criticality have been identified as resources for quantum enhanced sensitivity. In this section, we review various forms of quantum criticality for sensing and show that a common feature among all these systems is the emergence of a sort of “gap-closing” in their energy spectrum. This is an important observation as it clearly shows that any phase transition which is accompanied by gap-closing is a potential resource for quantum sensing.
@@
where $\eta$ is the scaling exponent of the long-range correlation function of the order parameter, as defined by $\langle\sigma_{i}^{x}\sigma_{i+r}^{x}\rangle\propto r^{2-d-\eta}$, at critical point for a $d$-dimensional model, and $d_{m}=\min(d,d_{c})$, with upper critical dimension $d_{c}=3$ for the transverse field Ising chain. The results, plotted in Fig. 3, attest that sub-shot-noise scaling is indeed achievable with such probes.
@@ IV.1 Free fermionic many-body probes
with $\bar{\circ}$ denoting complex conjugation and each column representing one eigenvector. The QFI for the ground state with respect to parameter $\theta$ can then be expressed as (Mukhopadhyay and Bayat, 2023)
@@ IV.2 Bosonic many-body probes
$H=-t\sum_{i}\left(b_{i}^{\dagger}b_{i+1}+b_{i+1}^{\dagger}b_{i}\right)-\mu\sum_{i}n_{i}+U\sum_{i}n_{i}\left(n_{i}-1\right)$ (48)
where $b_{i}(b_{i}^{\dagger})$-s are the Bosonic annihilation (creation) operators and $n_{i}{=}b_{i}^{\dagger}b_{i}$ is the number operator at site-$i$. The Bose-Hubbard can be implemented by placing cold atoms in Harmonic oscillator traps (Jaksch et al., 1998; Jaksch and Zoller, 2005; Lewenstein et al., 2007). At $T=0$, the phase diagram consists of a Mott insulating phase ($t\ll U$) and a Superfluid phase ($t\gg U$). The quantum phase transition between these phases from the Mott insulator phase to the Superfluid phase involves the spontaneous breaking of the $U(1)$ symmetry of the Bose-Hubbard Hamiltonian. Quantum sensing in these atomic systems combines the theoretical formalism of photon-interfermometry based metrology with reasonable ease of producing quantum states and robustness against particle loss. Considering a single-mode description of the resulting interacting BEC, where atom-atom interactions are described by a non-linear correction to the Harmonic trap potential, Gross et al obtained the first experimental demonstration of quantum-enhanced metrology in cold Bosonic atoms. Dunningham et al proposed the use of number-correlated squeezed BEC states (Dunningham et al., 2002), which are more robust to particle loss than NOON states. Another scheme by Dunningham and Burnett suggested that one can recreate the action of a (multiport) beam-splitter in an interferometer by simply tweaking the potential barrier between sites, thus enabling quantum-enhanced sensitivity (Dunningham and Burnett, 2004; Cooper et al., 2009). Heisenberg-limited quantum analogs of inertial guidance systems, also known as atomic gyroscopes, have been proposed in optical ring lattices (Cooper et al., 2010). While strictly not many-body systems, criticalities associated with superradiant phase transitions engineered in optical cavities with Rabi or Jaynes-Cummings type interactions have also received renewed attention for the possibility of quantum-enhanced sensing  Garbe et al. (2020). Moreover, effect of engineering anisotropies between coupling strengths of rotating wave and counter-rotating atom-cavity interactions of such models have also been studied in detail Wang et al. (2016); Zhu et al. (2023).
@@ IV.3 Spin chain probes with criticality-enhanced sensitivity
By factoring $J$ and defining $\theta{=}h/J$, the Hamiltonian takes the general form of Eq. (36). One can analytically solve the transverse Ising model Eq. (49) through Jordan-Wigner transformation (Pfeuty, 1970) which connects spin operators to fermionic ones as $c_{i}=\prod_{j=1}^{i-1}(-\sigma^{z}_{i})\sigma^{-}_{i}$ (and equivalently $c_{i}^{\dagger}=\prod_{j=1}^{i-1}(-\sigma^{z}_{i})\sigma^{+}_{i}$), where $\sigma^{\pm}{=}(\sigma^{x}\pm i\sigma^{y})/2$. The corresponding QFI for the dimensionless parameter $\theta$ has been analytically calculated for the ground state (Zanardi et al., 2008; Invernizzi et al., 2008) and has been shown to scale quadratically with system size, namely $I_{Q}\sim N^{2}$ near the critical point $h/J\rightarrow\pm 1$. Recently, a modular approach effectively augmenting these systems with periodic coupling defects (Mukhopadhyay and Bayat, 2023) has been proposed, which exploits the fact that multiple criticalities can be created in such a construction, all of which inherit the critical scaling behavior of the original chain. Alternatively, even for uniform chains, extending the ensemble measurement setup to real-time feedback control, i.e., adaptive strategies, allows one to overcome shot-noise scaling across the phase diagram  Salvia et al. (2023).
@@
Other quantum spin models have also been extensively studied in the literature for determination of quantum-enhanced scaling. In addition to their relevance for quantum sensing, such QFI-based analysis often reveals the nature of multipartite entanglement in many-body quantum systems. These include the 1D quantum XY models (Liu et al., 2013; Skotiniotis et al., 2015; Ye et al., 2016) with Dzyaloshinsky-Moriya interactions (Ozaydin and Altintas, 2015; Ye et al., 2018; Ozaydin and Altintas, 2020), XXZ models (Liu et al., 2015) which show Kosterlitz-Thouless phase transition (Zheng et al., 2015), and Heisenberg-type rotationally symmetric models (Ye et al., 2020; Lambert and Sørensen, 2019). See Ref. (Gu, 2010) for more references and earlier reviews. In addition to spinless fermion models, metrological advantage was also investigated for short-range spinful Fermi gases and QFI was shown to be an order parameter in such systems (Lucchesi and Chiofalo, 2019). Effect of lattice imperfections and disorder was considered in Ref. (Wei, 2019), whereby QFI analysis reveals that the two paradigmatic quantum disordered 1D lattices, namely Anderson (Anderson, 1958) and Aubry-Andre models (Aubry and André, 1980) belong to different universality classes. Experimental demonstration of quantum-enhanced critical sensing has also been recently performed (Liu et al., 2021) with a small NMR system. Extensions to multiparameter estimation by calculating the QFI matrix have also been considered (Bakmou et al., 2019; Di Fresco et al., 2022) along with experimental demonstration for NMR systems (Jiang et al., 2021c). Departure from $T=0$, i.e., ground state assumption has been considered in several directions. One is to consider metrology in finite-temperatures. Similar to other markers of quantumness, presence of a quantum critical region fanning out from the critical point at $T=0$ has been detected with QFI (Gabbrielli et al., 2018). Another popular direction of departure, namely. considering driven quantum many-body sensors instead of probes at equilibrium has been discussed in a later section of this review.
@@ IV.4 Quantum many-body probes with long-range interactions
Since local interactions are useful in quantum enhanced sensing, it is natural to ask whether criticality in models with longer-range interactions are also beneficial for metrology tasks. Unfortunately, exact results are rare for such systems and approximate methods, such as Density Matrix Renormalization Group (DMRG), are often unsatisfactory for generic Hamiltonians with long-range interactions because of an area law violation. However, QFI has indeed been studied in many systems showing interesting physics. Perhaps the most studied model is the Lipshin-Meshkov-Glick (LMG) Hamiltonian for N sites in 1D, which generalizes the anisotropic XY Hamiltonian by dropping the nearest-neighbour assumption. The Hamiltonian can be written as
@@
Garbe et al (Garbe et al., 2022) have recently considered the influence of various driving protocols on QFI-based sensing in LMG and other fully connected systems. In LMG and other fully connected models, all interactions are equally strong, in contrast to models where all interactions except the nearest or next nearest are ignored. Thus, one may wonder what happens for models where long-range interactions are present, but the interaction strength is not distance agnostic, e.g., it decays following a power law (Fernández-Lorenzo et al., 2018). Such models become extremely relevant for ion-trap based platforms, where the decay exponent can be fixed by tuning the lasers Monroe et al. (2021). In Ref. (Yang et al., 2022), the authors considered a long-range Kitaev chain, which is an extension of the $N$-site tight-binding model, and whose Hamiltonian is given by
$H_{\theta}$ $=-\frac{J}{2}\sum_{j=1}^{N}(c_{j}^{\dagger}c_{j+1}+c_{j+1}^{\dagger}c_{j})-\mu\sum_{j=1}^{N}(c_{j}^{\dagger}c_{j}-\frac{1}{2})$
@@
where $\kappa_{\eta,l}$ is the coefficient for interaction between two sites $l$ distance apart, $J$ is the nearest neighbour coupling and $\mu$ is the chemical potential. The power-law decay is a specific case when $\kappa_{\eta,\,l}\propto 1/l^{\eta}$. The exponent $\eta$ controls the range of interaction. For example, $\eta{=}0$ represents a fully connected graph in which all pairs interact equally while $\eta\rightarrow\infty$ represents nearest neighbor interaction. The following expression for QFI for estimating the parameter $\Delta$ was obtained in (Yang et al., 2022)

### v2

@@ II.4 Multi-parameter estimation
and $I(\boldsymbol{\theta})$ is the CFI matrix with elements Liu et al. (2019)
@@
One of the main differences between the single- and multi-parameter scenarios is that, while the CFI can be achieved in both scenarios, in principle, the QFI is attainable only in the single-parameter case and not for the jointly multi-parameter estimation Albarelli et al. (2020). This can be understood because, generally, there is no single symmetric logarithmic derivative $L_{i}$ that defines an optimal measurement basis suitable for all the unknown parameters. To quantify this notion of incompatibility between parameters Ragy et al. (2016)—those that cannot be estimated optimally with a single set of measurements—the incompatibility matrix $\boldsymbol{D}$ has been introduced Albarelli et al. (2020); Carollo et al. (2019a), also known as mean Uhlmann curvature Carollo et al. (2019a), with elements:
@@ II.5 Geometry of parameter estimation
Close links between Fisher information quantities and statistical distances have been found Wootters (1981); Braunstein and Caves (1994). In terms of the CFI, which quantifies the sensitivity of the probability distributions $p(x|\theta)$ concerning the unknown parameter $\theta$, the notion of distinguishability between different probability distributions among $M$ trials arises naturally (see Bengtsson and $\dot{Z}$yczkowski Lehmann and Casella (1998), and Kok and Lovett Kok and Lovett (2010)). It turns out that distinguishing between two probabilities $p(x|\theta_{1})$ and $p(x|\theta_{2})$, taking $\text{d}s$ ($\text{d}s$ being the infinitesimal distance on the probability simplex space) along a line element $\text{d}\theta$, results in Lehmann and Casella (1998); Kok and Lovett (2010)
@@
which is the CFI of Eq. (3). Further analysis leads to:
@@
where $\delta\theta$ is the segment of the path in the probability simplex—rather than the variance of $\tilde{\theta}$. Note that the above has been derived solely from a distance measure (a metric) by distinguishing probability distributions on the simplex space. It has also been shown that for distinguishing two probability distributions, one can use the relative entropy (not a metric) for this purpose, and consequently, a close connection between the CFI matrix and the relative entropy emerges Sidhu and Kok (2020).
@@
For pure states $\rho_{1}=|\psi_{1}\rangle\langle\psi_{1}|$ and $\rho_{2}=|\psi_{2}\rangle\langle\psi_{2}|$, the Uhlmann fidelity reduces to $F(\rho_{1},\rho_{2})=|\langle\psi_{1}|\psi_{2}\rangle|$.
@@ II.6 Alternatives to QFI for Metrology
where $I(\theta)$ is the QFI for $\rho(\theta)$ with respect to the parameter $\theta$. One can note that this bound reduces to the biased version of the Bayesian Quantum Cramér-Rao bound bound for a uniform prior. Secondly, and of significant relevance to many-body sensing, the QFI is not fully expressible in terms of experimentally easily accessible one and two-point correlation functions(Gabbrielli et al., 2018). This was addressed recently in Ref. (Lepori et al., 2023), where an alternative functional dependent only on one and two-point correlators was proposed. More economic sensing with fewer copies was studied in Ref. Zhang and Tong (2024).
@@ III Interferometric Quantum Sensing
Note that the quantum enhanced sensitivity with interferometric approach relies on creating GHZ-type entangled states or equivalently $N00N$ Mitchell et al. (2004) or squeezed states Ma et al. (2011); Maccone and Riccardi (2020). Preparation of these special forms of entangled states is not an easy task and is resource consuming. In fact, delocalization of quantum correlation is found to be essential for fast generation of metrologically useful entangled states Chu et al. (2024). In Ref. Chu et al. (2023), a strong metrological limit for the preparation time of such large-scale entanglement has been established through studying the Lieb-Robinson light cone.
@@ IV Equilibrium Quantum Sensing: Second Order Quantum Phase Transition
An alternative method to interferometric sensing is to harness the interaction between the particles in strongly correlated many-body systems. The interplay between different parts of the Hamiltonian can induce phase transition in the system, also known as quantum criticality. Many-body systems exhibiting quantum criticality have been identified as resources for quantum enhanced sensitivity. Recently, C. Hotter et al. proposed a unified protocol that combines critical and conventional approaches to metrology Hotter et al. (2024). A more fundamental limit to many-body quantum metrology was recently obtained in Ref. Abiuso et al. (2024). In this section, we review various forms of quantum criticality for sensing and show that a common feature among all these systems is the emergence of a sort of “gap-closing” in their energy spectrum. This is an important observation as it clearly shows that any phase transition which is accompanied by gap-closing is a potential resource for quantum sensing.
@@
where $\eta$ is the scaling exponent of the long-range correlation function of the order parameter, as defined by $\langle\sigma_{i}^{x}\sigma_{i+r}^{x}\rangle\propto r^{2-d-\eta}$, at critical point for a $d$-dimensional model, and $d_{m}=\min(d,d_{c})$, with upper critical dimension $d_{c}=3$ for the transverse field Ising chain. The results, plotted in Fig. 3, attest that sub-shot-noise scaling is indeed achievable with such probes. Similar results have also been found in Ref. Li et al. (2024) in which interferometric quantum enhanced sensitivity, using the critical ground state of a many-body system, can be exploited for detecting quantum phase transitions.
@@ IV.1 Free fermionic many-body probes
with $\bar{\circ}$ denoting complex conjugation and each column representing one eigenvector. The QFI for the ground state with respect to parameter $\theta$ can then be expressed as (Mukhopadhyay and Bayat, 2024)
@@ IV.2 Bosonic many-body probes
$H=-t\sum_{i}\left(b_{i}^{\dagger}b_{i+1}+b_{i+1}^{\dagger}b_{i}\right)-\mu\sum_{i}n_{i}+\frac{U}{2}\sum_{i}n_{i}\left(n_{i}-1\right)$ (48)
where $b_{i}(b_{i}^{\dagger})$-s are the Bosonic annihilation (creation) operators, $n_{i}{=}b_{i}^{\dagger}b_{i}$ is the number operator at site $i$, and $t,U,\mu$ are the hopping, onsite interaction, and chemical potential paremeters, respectively. The Bose-Hubbard model can be implemented by loading cold bosonic atoms in Harmonic oscillator traps (Jaksch et al., 1998; Jaksch and Zoller, 2005; Lewenstein et al., 2007). At $T=0$, the phase diagram consists of a Mott insulating phase ($t\ll U$) and a Superfluid phase ($t\gg U$). The quantum phase transition between these phases from the Mott insulator phase to the Superfluid phase involves the spontaneous breaking of the $U(1)$ symmetry of the Bose-Hubbard Hamiltonian. Quantum sensing in these atomic systems combines the theoretical formalism of photon-interfermometry based metrology with reasonable ease of producing quantum states and robustness against particle loss. Considering a single-mode description of the resulting interacting BEC, where atom-atom interactions are described by a non-linear correction to the Harmonic trap potential, Gross et al obtained the first experimental demonstration of quantum-enhanced metrology in cold Bosonic atoms Gross et al. (2010). Dunningham et al proposed the use of number-correlated squeezed BEC states (Dunningham et al., 2002), which are more robust to particle loss than NOON states. Another scheme by Dunningham and Burnett suggested that one can recreate the action of a (multiport) beam-splitter in an interferometer by simply tweaking the potential barrier between sites, thus enabling quantum-enhanced sensitivity (Dunningham and Burnett, 2004; Cooper et al., 2009). Heisenberg-limited quantum analogs of inertial guidance systems, also known as atomic gyroscopes, have been proposed in optical ring lattices (Cooper et al., 2010). A multiconfigurational Hartree type self-consistent theoretical method for calculation of quantum metrological quantities for interacting trapped Bosonic systems has also been proposed Baak and Fischer (2024) . Precision estimation of chemical potential has also been studied in Bosonic quantum gases Marzolino and Braun (2013). While strictly not many-body systems, criticalities associated with superradiant phase transitions engineered in optical cavities with Rabi or Jaynes-Cummings type interactions have also received renewed attention for the possibility of quantum-enhanced sensing  Garbe et al. (2020). Moreover, effect of engineering anisotropies between coupling strengths of rotating wave and counter-rotating atom-cavity interactions of such models have also been studied in detail Wang et al. (2016); Zhu et al. (2023).
@@ IV.3 Spin chain probes with criticality-enhanced sensitivity
By factoring $J$ and defining $\theta{=}h/J$, the Hamiltonian takes the general form of Eq. (36). One can analytically solve the transverse Ising model Eq. (49) through Jordan-Wigner transformation (Pfeuty, 1970) which connects spin operators to fermionic ones as $c_{i}=\prod_{j=1}^{i-1}(-\sigma^{z}_{i})\sigma^{-}_{i}$ (and equivalently $c_{i}^{\dagger}=\prod_{j=1}^{i-1}(-\sigma^{z}_{i})\sigma^{+}_{i}$), where $\sigma^{\pm}{=}(\sigma^{x}\pm i\sigma^{y})/2$. The corresponding QFI for the dimensionless parameter $\theta$ has been analytically calculated for the ground state (Zanardi et al., 2008; Invernizzi et al., 2008) and has been shown to scale quadratically with system size, namely $I_{Q}\sim N^{2}$ near the critical point $h/J\rightarrow\pm 1$. Recently, a modular approach effectively augmenting these systems with periodic coupling defects (Mukhopadhyay and Bayat, 2024) has been proposed, which exploits the fact that multiple criticalities can be created in such a construction, all of which inherit the critical scaling behavior of the original chain. Alternatively, even for uniform chains, extending the ensemble measurement setup to real-time feedback control, i.e., adaptive strategies, allows one to overcome shot-noise scaling across the phase diagram  Salvia et al. (2023).
@@
Other quantum spin models have also been extensively studied in the literature for determination of quantum-enhanced scaling. In addition to their relevance for quantum sensing, such QFI-based analysis often reveals the nature of multipartite entanglement in many-body quantum systems. These include the 1D quantum XY models (Liu et al., 2013; Skotiniotis et al., 2015; Ye et al., 2016) with Dzyaloshinsky-Moriya interactions (Ozaydin and Altintas, 2015; Ye et al., 2018; Ozaydin and Altintas, 2020), XXZ models (Liu et al., 2015) which show Kosterlitz-Thouless phase transition (Zheng et al., 2015), Heisenberg-type rotationally symmetric models (Ye et al., 2020; Lambert and Sørensen, 2019), and even applied to the quantum Hall probes Giraud et al. (2024). See Ref. (Gu, 2010) for more references and earlier reviews. In addition to spinless fermion models, metrological advantage was also investigated for short-range spinful Fermi gases and QFI was shown to be an order parameter in such systems (Lucchesi and Chiofalo, 2019). Higher dimensional spin chains have also been shown to improve magnetometry with arbitrary spin-number transverse Ising models Singh et al. (2024). Effect of lattice imperfections and disorder was considered in Ref. (Wei, 2019), whereby QFI analysis reveals that the two paradigmatic quantum disordered 1D lattices, namely Anderson  (Anderson, 1958) and Aubry-Andre models (Aubry and André, 1980) belong to different universality classes. Experimental demonstration of quantum-enhanced critical sensing has also been recently performed (Liu et al., 2021) with a small NMR system. Extensions to multiparameter estimation by calculating the QFI matrix have also been considered (Bakmou et al., 2019; Di Fresco et al., 2022) along with experimental demonstration for NMR systems (Jiang et al., 2021c). Departure from $T=0$, i.e., ground state assumption has been considered in several directions. One direction is to consider metrology in finite-temperatures. Similar to other markers of quantumness, presence of a quantum critical region fanning out from the critical point at $T=0$ has been detected with QFI (Gabbrielli et al., 2018). Another popular direction of departure, namely considering driven quantum many-body sensors instead of probes at equilibrium has been discussed in a later section of this review.
@@ IV.4 Quantum many-body probes with long-range interactions
Since local interactions are useful in quantum enhanced sensing, it is natural to ask whether criticality in models with longer-range interactions are also beneficial for metrology tasks Monika et al. (2023). Unfortunately, exact results are rare for such systems and approximate methods, such as Density Matrix Renormalization Group (DMRG), are often unsatisfactory for generic Hamiltonians with long-range interactions because of an area law violation. However, QFI has indeed been studied in many systems showing interesting physics. Perhaps the most studied model is the Lipshin-Meshkov-Glick (LMG) Hamiltonian for N sites in 1D, which generalizes the anisotropic XY Hamiltonian by dropping the nearest-neighbour assumption. The Hamiltonian can be written as
@@
Garbe et al (Garbe et al., 2022) have recently considered the influence of various driving protocols on QFI-based sensing in LMG and other fully connected systems. In LMG and other fully connected models, all interactions are equally strong, in contrast to models where all interactions except the nearest or next nearest are ignored. Thus, one may wonder what happens for models where long-range interactions are present, but the interaction strength is not distance agnostic, e.g., it decays following a power law (Fernández-Lorenzo et al., 2018). Such models become extremely relevant for ion-trap based platforms, where the decay exponent can be fixed by tuning the lasers Monroe et al. (2021). In Ref. (Yang et al., 2022a), the authors considered a long-range Kitaev chain, which is an extension of the $N$-site tight-binding model, and whose Hamiltonian is given by
$H$ $=-\frac{J}{2}\sum_{j=1}^{N}(c_{j}^{\dagger}c_{j+1}+c_{j+1}^{\dagger}c_{j})-\mu\sum_{j=1}^{N}(c_{j}^{\dagger}c_{j}-\frac{1}{2})$
@@
where $\kappa_{\eta,l}$ is the coefficient for interaction between two sites $l$ distance apart, $J$ is the nearest neighbour coupling and $\mu$ is the chemical potential. The power-law decay is a specific case when $\kappa_{\eta,\,l}\propto 1/l^{\eta}$. The exponent $\eta$ controls the range of interaction. For example, $\eta{=}0$ represents a fully connected graph in which all pairs interact equally while $\eta\rightarrow\infty$ represents nearest neighbor interaction. The following expression for QFI for estimating the parameter $\Delta$ was obtained in (Yang et al., 2022a)